# Meta-Learning Driven Movable-Antenna-assisted Full-Duplex RSMA for Multi-User Communication: Performance and Optimization

### v2

@@ II-A Transmission Model
where $\textbf{W}_{c}\in\mathbb{C}^{N_{T}\times 1}$ and $\textbf{W}_{p,d}\in\mathbb{C}^{N_{T}\times 1}$ represent the transmitting beamforming vectors to transmit a common stream for all DL UEs and a private stream for DL UE-$d$, respectively. We assume a FD BS system can transmit and receive signals at the same time. In other words, in the same resource block, the BS transmits a signal to the DL UEs and receives a signal from UL UEs. In particular, utilizing the UL RSMA principle, the messages of UL UEs are split into two sub-messages except for one UL UE. More specifically, we split the messages of $U-1$ UEs into two sub-messages, and the last UE’s message is kept without splitting. Because it has been shown in [20] that for a RSMA system with $U$ UL UEs, splitting the messages $U-1$ is sufficient to achieve the capacity region. Hence, we split each UL UE’s message into two sub-messages; the set of sub-messages is denoted as $J\in[1,2]$. Therefore, the transmitted signal of UE-$u,\forall u\in\mathcal{U}$ to the BS can be given by,
@@
$y_{d}=\underbrace{\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})(\textbf{W}_{c}s_{c}+\textbf{W}_{p,d}s_{p,d})}_{\textrm{desired signal}}+\underbrace{\sum_{d^{\prime}\in\mathcal{D},d^{\prime}\neq d}\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{p^{\prime},d^{\prime}}s_{p^{\prime},d^{\prime}}}_{\textrm{interference from private streams of other DLUEs}}+\underbrace{\sum_{u\in\mathcal{U}}h_{u,d}(\textbf{t}^{UL},\textbf{r}^{DL})x_{u}}_{\textrm{intra-cell interference from UL UEs}}+n_{d},$ (4)
@@
$\hat{\textbf{y}}=\underbrace{\sum_{u\in\mathcal{U}}\textbf{h}_{u}(\textbf{t}^{UL},\textbf{r}_{n_{r}}^{BS})x_{u}}_{\textrm{desired signal}}+\underbrace{\textbf{H}_{SI}(\textbf{t}_{n_{t}}^{BS},\textbf{r}_{n_{r}}^{BS})\hat{\textbf{x}}}_{\textrm{self-interference}}+\textbf{n}_{BS},$ (5)
@@ II-B1 Achievable rate for DL
$\footnotesize R_{c,d}=\log_{2}\left(1+\frac{|\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{c}|^{2}}{\sum_{d\in\mathcal{D}}|\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{p,d}|^{2}+\sum_{u=1}^{U-1}(\sum_{j\in J}P_{u,j}|h_{u,d}(\textbf{t}^{UL},\textbf{r}^{DL})|^{2})+P_{U}|h_{U,d}(\textbf{t}^{UL},\textbf{r}^{DL})|^{2}+\sigma_{d}^{2}}\right).$ (6)
@@
$\footnotesize R_{p,d}=\log_{2}\left(1+\frac{|\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{p,d}|^{2}}{\sum_{d^{\prime}\in\mathcal{D},d^{\prime}\neq d}|\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{p^{\prime},d^{\prime}}|^{2}+\sum_{u=1}^{U-1}(\sum_{j\in J}P_{u,j}|h_{u,d}(\textbf{t}^{UL},\textbf{r}^{DL})|^{2})+P_{U}|h_{U,d}(\textbf{t}^{UL},\textbf{r}^{DL})|^{2}+\sigma_{d}^{2}}\right).$ (7)
@@
where $C_{d}$ represents the fraction of $R_{c}$ assigned to UE-$d$, $\forall d\in\mathcal{D}$.
@@ II-B2 Achievable rate analysis for UL
It is worth mentioning that in the UL scenario, decoding order plays a crucial role in processing the received signals. Let the decoding order of sub-messages and messages at the BS be represented by the set $\pi=[\pi_{u,j},\pi_{U}|\forall j\in J,\forall u\in\mathcal{U}]$ where the elements are arranged in ascending order. Here, $\pi_{u,j}$ denotes the decoding order of the sub-message $s_{u,j}$, while $\pi_{U}$ corresponds to the decoding orders of the sub-message $s_{U}$. Specifically, sub-message $s_{u,j}$ is decoded first if its decoding order satisfies $\pi_{u,j}<\pi_{u^{\prime},j^{\prime}}$ or $\pi_{u,j}<\pi_{U}$ where $u^{\prime}\neq u$ and $j\neq j^{\prime}$, while treating all remaining sub-messages/message as interference. Once $s_{u,j}$ has been successfully decoded, it is removed from the received signal through SIC. The process then continues with the next sub-message or message according to the predefined decoding order. The adopted decoding order for this work is specified in simulation section. Consequently, the achievable rate to decode sub-message $s_{u,j}$ at BS can be denoted as (10),
$\footnotesize R_{u,j}=\log_{2}\left(1+\frac{P_{u,j}|\textbf{Z}_{u,j}^{H}\textbf{h}_{u}(\textbf{t}^{UL},\textbf{r}_{n_{r}}^{BS})|^{2}}{{}_{j^{\prime}\neq j}P_{u,j^{\prime}}|\textbf{Z}_{u,j}^{H}\textbf{h}_{u}(\textbf{t}^{UL},\textbf{r}_{nr}^{BS})|^{2}+I+\sigma_{u}^{2}||\textbf{Z}_{u,j}^{H}||_{2}^{2}}\right).$ (10)
@@
$\footnotesize\begin{split}&I=\sum_{u^{\prime}\neq u}^{u^{\prime}\in(U-1)}(\sum_{j\in J}P_{u^{\prime},j}|\textbf{Z}_{u,j}^{H}\textbf{h}_{u^{\prime}}(\textbf{t}^{UL},\textbf{r}_{nr}^{BS})|^{2})+P_{U}|\textbf{Z}_{u,j}^{H}\textbf{h}_{U}(\textbf{t}^{UL},\textbf{r}_{nr}^{BS})|^{2}+\\ &|\textbf{Z}_{u,j}^{H}|^{2}|\textbf{H}_{SI}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}_{n_{r}}^{BS})(\textbf{W}_{c}+\sum_{d\in\mathcal{D}}\textbf{W}_{p,d})|^{2}.\end{split}$
@@
$R_{U}=\log_{2}\left(1+\frac{P_{U}|\textbf{Z}_{U}^{H}\textbf{h}_{U}(\textbf{t}^{UL},\textbf{r}_{n_{r}}^{BS})|^{2}}{\sum_{u^{\prime}\in\mathcal{U},u^{\prime}\neq U}(\sum_{j\in J}P_{u^{\prime},j}|\textbf{Z}_{U}^{H}\textbf{h}_{u^{\prime}}(\textbf{t}^{UL},\textbf{r}_{n_{r}}^{BS})|^{2})+|\textbf{Z}_{U}^{H}|^{2}|\textbf{H}_{SI}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}_{n_{r}}^{BS})(\textbf{W}_{c}+\sum_{d\in\mathcal{D}}\textbf{W}_{p,d})|^{2}+\sigma_{U}^{2}||\textbf{Z}_{U}^{H}||^{2}_{2}}\right).$ (11)
@@ II-C Field-Response Based Channel Model
Following [3], we model all the communication channels based on the field-response channel model. For MA systems, the channel response is influenced by both the propagation environment and the positioning of the antennas. Similar to [21], we assume that the antenna’s movement region is significantly smaller than the propagation distance between the transmitter and receiver, ensuring that the far-field condition holds at both ends. This assumption is reasonable because the typical size of the antenna moving region is in the order of several to tens of wavelengths. As a result, the plane-wave model can be applied to characterize the field response between the transmit and receive regions. This implies that the angles of departure (AoD)s, angles of arrival (AoA)s, and the amplitudes of the complex coefficients for multiple channel paths remain unchanged regardless of the MAs’ positions. In contrast, only the phases of the multi-path channels vary within the transmit and receive regions.
@@ II-C1 DL Channels
At the BS side, we denote the number of transmit paths as $L^{BS}_{t}$ from the BS. We assume that the elevation and azimuth AoDs for the $j$-th transmit path between BS and UE-$d$ can be denoted as $\theta_{t,j}^{BS}\in[-\pi/2,\pi/2]$ and $\phi_{t,j}^{BS}\in[-\pi/2,\pi/2],1\leq j\leq L_{t}^{DL}$. At the UE side, we denote the number of receive paths as $L_{r}^{d}$. The elevation and azimuth AoAs of the $i$-th receive paths are respectively denoted as $\theta_{r,i}^{d}\in[-\pi/2,\pi/2]$, $\phi_{r,i}^{d}\in[-\pi/2,\pi/2]$, $1\leq i\leq L_{r}^{d}$. We define the path-response matrix (PRM) $\boldsymbol{\Sigma}_{d}\in L_{r}^{d}\times L^{BS}_{t}$ to represent the response from the transmit and reference position $\textbf{t}_{n_{t}}^{BS}(0)=[0,0]$ to the receive reference position $\textbf{r}^{DL}(0)=[0,0]$. Now, the difference between position $\textbf{r}^{DL}=[x_{r},y_{r}]$ and the reference point $\textbf{r}^{DL}[0]$ for the $i$-th receive path, $1\leq i\leq L_{r}^{d}$ is given as follows,
@@
$\footnotesize\textbf{f}_{d}({\textbf{r}^{DL}})=[e^{j\frac{2\pi}{\lambda}\rho_{r,1}(\textbf{r}^{DL})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{r,L_{r}^{d}}(\textbf{r}^{DL})}]^{T}\in\mathbb{C}^{L_{r}^{d}\times 1}.$ (14)
@@
$\footnotesize\textbf{g}(\textbf{t}_{n_{t}}^{BS})=[e^{j\frac{2\pi}{\lambda}\rho_{t,1}(\textbf{t}_{n_{t}}^{BS})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{t,L^{BS}_{t}}(\textbf{t}_{n_{t}}^{BS})}]^{T}\in\mathbb{C}^{{L_{t}^{BS}}\times 1},$ (15)
@@ II-C2 UL channels
$\footnotesize\textbf{r}_{n_{r}}^{BS}=[e^{j\frac{2\pi}{\lambda}\rho_{r,1}(\textbf{r}_{n_{r}}^{BS})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{r,L^{BS}_{r}}(\textbf{r}_{n_{r}}^{BS})}]^{T}\in\mathbb{C}^{{L}_{r}^{BS}\times 1}$ (18)
@@
$\footnotesize\textbf{g}_{u}(\textbf{t}^{UL})=[e^{j\frac{2\pi}{\lambda}\rho_{t,1}(\textbf{t}^{UL})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{t,L^{d}_{t}}(\textbf{t}^{UL})}]^{T}\in\mathbb{C}^{L_{t}^{d}\times 1}$ (19)
@@ II-C3 SI channel
$\footnotesize\textbf{f}_{SI}(\textbf{r}_{nr}^{BS})=[e^{j\frac{2\pi}{\lambda}\rho_{r,1}(\textbf{r}_{n_{r}}^{BS})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{r,L^{SI}_{r}}(\textbf{r}_{n_{r}})}]^{T}\in\mathbb{C}^{L_{r}^{SI}\times 1},$ (22)
$\footnotesize\boldsymbol{\Sigma}_{SI}\in\mathbb{C}^{L_{r}^{SI}\times L_{t}^{SI}},$ (23)
$\footnotesize\textbf{G}_{SI}=[\textbf{g}_{SI}(t_{1}^{BS}),\dots,\textbf{g}_{SI}(t_{N_{T}}^{BS})]\in\mathbb{C}^{L_{t}^{SI}\times N_{T}},$ (24)
$\footnotesize\textbf{g}_{SI}(\textbf{t}_{n_{t}}^{BS})=[e^{j\frac{2\pi}{\lambda}\rho_{n_{t},1}(\textbf{t}_{nt}^{BS})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{n_{t},L^{SI}_{t}}(\textbf{t}_{n_{t}}^{BS})}]^{T}\in\mathbb{C}^{L_{t}^{SI}\times 1}.$ (25)
@@ II-C4 Interference channels
$\footnotesize\textbf{f}_{u}(\textbf{r}^{DL})=[e^{j\frac{2\pi}{\lambda}\rho_{r,1}(\textbf{r}^{DL})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{r,L^{ud}_{r}}(\textbf{r}^{DL})}]^{T}\in\mathbb{C}^{L_{r}^{ud}\times 1},$ (27)
@@
$\footnotesize\textbf{g}_{u}(\textbf{t}^{UL})=[e^{j\frac{2\pi}{\lambda}\rho_{t,1}(\textbf{t}^{UL})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{t,L^{ud}_{t}}(\textbf{t}^{UL})}]^{T}\in\mathbb{C}^{L_{t}^{ud}\times 1}.$ (29)

### v3

@@ II-A Transmission Model
where $\textbf{W}_{c}\in\mathbb{C}^{N_{T}\times 1}$ and $\textbf{W}_{p,d}\in\mathbb{C}^{N_{T}\times 1}$ represent the transmitting beamforming vectors to transmit a common stream for all DL UEs and a private stream for DL UE-$d$, respectively. We assume a FD BS system can transmit and receive signals at the same time. In other words, in the same resource block, the BS transmits a signal to the DL UEs and receives a signal from UL UEs. In particular, utilizing the UL RSMA principle, the messages of UL UEs are split into two sub-messages except for one UL UE. More specifically, we split the messages of U1U-1 UEs into two sub-messages, and the last UE’s message is kept without splitting. Because it has been shown in [liu2020rate] that for a RSMA system with $U$ UL UEs, splitting the messages U1U-1 is sufficient to achieve the capacity region. Hence, we split each UL UE’s message into two sub-messages; the set of sub-messages is denoted as $J\in[1,2]$. Therefore, the transmitted signal of UE-$u,\forall u\in\mathcal{U}$ to the BS can be given by,
@@
[…]
@@
$\hat{\textbf{y}}=\underbrace{\sum_{u\in\mathcal{U}}\textbf{h}_{u}(\textbf{t}^{UL},\textbf{r}_{n_{r}}^{BS})x_{u}}_{\textrm{desired signal}}+\underbrace{\textbf{H}^{H}_{SI}(\textbf{t}_{n_{t}}^{BS},\textbf{r}_{n_{r}}^{BS})\hat{\textbf{x}}}_{\textrm{self-interference}}+\textbf{n}_{BS},$ (5)
@@ II-B1 Achievable rate for DL
$R_{c,d}=\log_{2}\left(1+\frac{|\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{c}|^{2}}{\sum_{d\in\mathcal{D}}|\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{p,d}|^{2}+\sum_{u=1}^{U-1}(\sum_{j\in J}P_{u,j}|h_{u,d}(\textbf{t}^{UL},\textbf{r}^{DL})|^{2})+P_{U}|h_{U,d}(\textbf{t}^{UL},\textbf{r}^{DL})|^{2}+\sigma_{d}^{2}}\right).$ (6)
@@
$R_{p,d}=\log_{2}\left(1+\frac{|\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{p,d}|^{2}}{\sum_{d^{\prime}\in\mathcal{D},d^{\prime}\neq d}|\textbf{h}_{d}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}^{DL})\textbf{W}_{p^{\prime},d^{\prime}}|^{2}+\sum_{u=1}^{U-1}(\sum_{j\in J}P_{u,j}|h_{u,d}(\textbf{t}^{UL},\textbf{r}^{DL})|^{2})+P_{U}|h_{U,d}(\textbf{t}^{UL},\textbf{r}^{DL})|^{2}+\sigma_{d}^{2}}\right).$ (7)
@@
where $C_{d}$ represents the fraction of $R_{c}$ assigned to UE-$d$.
@@ II-B2 Achievable rate analysis for UL
It is worth mentioning that in the UL scenario, decoding order plays a crucial role in processing the received signals. Let the decoding order of sub-messages and messages at the BS be represented by the set $\pi=[\pi_{u,j},\pi_{U}|\forall j\in J,\forall u\in\mathcal{U}]$ where the elements are arranged in ascending order. Here, $\pi_{u,j}$ denotes the decoding order of the sub-message $s_{u,j}$, while $\pi_{U}$ corresponds to the decoding orders of the sub-message $s_{U}$. Specifically, sub-message $s_{u,j}$ is decoded first if its decoding order satisfies $\pi_{u,j}<\pi_{u^{\prime},j^{\prime}}$ or $\pi_{u,j}<\pi_{U}$ where $u^{\prime}\neq u$ and $j\neq j^{\prime}$, while treating all remaining sub-messages/message as interference. Once $s_{u,j}$ has been successfully decoded, it is removed from the received signal through SIC. The process then continues with the next sub-message or message according to the predefined decoding order. The adopted decoding order for this work is specified in the simulation section. Consequently, the achievable rate to decode sub-message $s_{u,j}$ at BS can be denoted as (10),
[…]
@@
$\begin{split}&I=\sum_{u^{\prime}\neq u}^{u^{\prime}\in(U-1)}(\sum_{j\in J}P_{u^{\prime},j}|\textbf{Z}_{u,j}^{H}\textbf{h}_{u^{\prime}}(\textbf{t}^{UL},\textbf{r}_{nr}^{BS})|^{2})+\\ &P_{U}|\textbf{Z}_{u,j}^{H}\textbf{h}_{U}(\textbf{t}^{UL},\textbf{r}_{nr}^{BS})|^{2}+|\textbf{Z}_{u,j}^{H}|^{2}|\textbf{H}_{SI}^{H}(\textbf{t}_{n_{t}}^{BS},\textbf{r}_{n_{r}}^{BS})(\textbf{W}_{c}+\sum_{d\in\mathcal{D}}\textbf{W}_{p,d})|^{2}.\end{split}$
@@
[…]
@@ II-C Field-Response Based Channel Model
Following [ding2024movable], we model all the communication channels based on the field-response channel model. For MA systems, the channel response is influenced by both the propagation environment and the positioning of the antennas. Similar to [ma2023mimo], we assume that the antenna’s movement region is significantly smaller than the propagation distance between the transmitter and receiver, ensuring that the far-field condition holds at both ends. This assumption is reasonable because the typical size of the antenna moving region is in the order of several to tens of wavelengths. As a result, the plane-wave model can be applied to characterize the field response between the transmit and receive regions. This implies that the angles of departure (AoD)s, angles of arrival (AoA)s, and the amplitudes of the complex coefficients for multiple channel paths remain unchanged regardless of the MAs’ positions. In contrast, only the phases of the multi-path channels vary within the transmit and receive regions.
@@ II-C1 DL Channels
At the BS side, we denote the number of transmit paths as $L^{BS}_{t}$ from the BS. We assume that the elevation and azimuth AoDs for the $j$-th transmit path between BS and UE-$d$ can be denoted as $\theta_{t,j}^{BS}\in[-\pi/2,\pi/2]$ and $\phi_{t,j}^{BS}\in[-\pi/2,\pi/2],1\leq j\leq L_{t}^{DL}$. At the UE side, we denote the number of receive paths as $L_{r}^{d}$. The elevation and azimuth AoAs of the $i$-th receive paths are respectively denoted as $\theta_{r,i}^{d}\in[-\pi/2,\pi/2]$, $\phi_{r,i}^{d}\in[-\pi/2,\pi/2]$, $1\leq i\leq L_{r}^{d}$. We define the path-response matrix (PRM) $\boldsymbol{\Sigma}_{d}\in L_{r}^{d}\times L^{BS}_{t}$ to represent the response from the transmit and reference position $\textbf{t}_{n_{t}}^{BS}(0)=[0,0]$ to the receive reference position $\textbf{r}^{DL}(0)=[0,0]$. Now, the difference between position $\textbf{r}^{DL}=[x_{r},y_{r}]$ and the reference point $\textbf{r}^{DL}[0]$ for the $i$-th receive path, $1\leq i\leq L_{r}^{d}$ is given as follows:
@@
$\textbf{f}_{d}({\textbf{r}^{DL}})=[e^{j\frac{2\pi}{\lambda}\rho_{r,1}(\textbf{r}^{DL})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{r,L_{r}^{d}}(\textbf{r}^{DL})}]^{T}\in\mathbb{C}^{L_{r}^{d}\times 1}.$ (14)
@@
$\textbf{g}(\textbf{t}_{n_{t}}^{BS})=[e^{j\frac{2\pi}{\lambda}\rho_{t,1}(\textbf{t}_{n_{t}}^{BS})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{t,L^{BS}_{t}}(\textbf{t}_{n_{t}}^{BS})}]^{T}\in\mathbb{C}^{{L_{t}^{BS}}\times 1},$ (15)
@@ II-C2 UL channels
$\textbf{f}_{b}(\textbf{r}_{n_{r}}^{BS})=[e^{j\frac{2\pi}{\lambda}\rho_{r,1}(\textbf{r}_{n_{r}}^{BS})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{r,L^{BS}_{r}}(\textbf{r}_{n_{r}}^{BS})}]^{T}\in\mathbb{C}^{{L}_{r}^{BS}\times 1}$ (18)
@@
$\textbf{g}_{u}(\textbf{t}^{UL})=[e^{j\frac{2\pi}{\lambda}\rho_{t,1}(\textbf{t}^{UL})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{t,L^{d}_{t}}(\textbf{t}^{UL})}]^{T}\in\mathbb{C}^{L_{t}^{d}\times 1}$ (19)
@@ II-C3 SI channel
$\textbf{f}_{SI}(\textbf{r}_{nr}^{BS})=[e^{j\frac{2\pi}{\lambda}\rho_{r,1}(\textbf{r}_{n_{r}}^{BS})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{r,L^{SI}_{r}}(\textbf{r}_{n_{r}})}]^{T}\in\mathbb{C}^{L_{r}^{SI}\times 1},$ (22)
$\boldsymbol{\Sigma}_{SI}\in\mathbb{C}^{L_{r}^{SI}\times L_{t}^{SI}},$ (23)
$\textbf{G}_{SI}=[\textbf{g}_{SI}(t_{1}^{BS}),\dots,\textbf{g}_{SI}(t_{N_{T}}^{BS})]\in\mathbb{C}^{L_{t}^{SI}\times N_{T}},$ (24)
$\textbf{g}_{SI}(\textbf{t}_{n_{t}}^{BS})=[e^{j\frac{2\pi}{\lambda}\rho_{n_{t},1}(\textbf{t}_{nt}^{BS})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{n_{t},L^{SI}_{t}}(\textbf{t}_{n_{t}}^{BS})}]^{T}\in\mathbb{C}^{L_{t}^{SI}\times 1}.$ (25)
@@ II-C4 Interference channels
$\textbf{f}_{u}(\textbf{r}^{DL})=[e^{j\frac{2\pi}{\lambda}\rho_{r,1}(\textbf{r}^{DL})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{r,L^{ud}_{r}}(\textbf{r}^{DL})}]^{T}\in\mathbb{C}^{L_{r}^{ud}\times 1},$ (27)
@@
$\textbf{g}_{u}(\textbf{t}^{UL})=[e^{j\frac{2\pi}{\lambda}\rho_{t,1}(\textbf{t}^{UL})},\dots,e^{j\frac{2\pi}{\lambda}\rho_{t,L^{ud}_{t}}(\textbf{t}^{UL})}]^{T}\in\mathbb{C}^{L_{t}^{ud}\times 1}.$ (29)